The sum of two numbers is 16. Their difference is 40. What are the numbers?
The sum of three consecutive integers is 96. What are the numbers?

Answer by Alan3354(69443)

The sum of two numbers is 16. Their difference is 40. What are the numbers?
16/2 = 8
40/2 = 20
--> 8 ± 20
= -12 & 28
-------------
The sum of three consecutive integers is 96. What are the numbers?
96/3 = 32, the middle number